OpenStudy (anonymous):

LAST ONE..for now ;) Brian is creating a collage on a piece of cardboard that has an area of 160r4 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the seventh power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

hartnn (hartnn):

total area:160r^4 area of one piece = \[\sqrt{r^7}\] so total pieces = \[\frac{160 r^4}{\sqrt{r^7}}\] can u solve further?

OpenStudy (anonymous):

no? im so confused..:(

hartnn (hartnn):

which step did u not understand? or u confused with simplification??

OpenStudy (anonymous):

yea i realy just dont get the whole concept of this..:p so you divide them but like how?

hartnn (hartnn):

ok, tell me if i have 1 big rectangle with area=100,how many rectangles of area 10 would i fit in??

OpenStudy (anonymous):

10

hartnn (hartnn):

what did u do?? divide,right?? this is the same thing...thats why i divided...

OpenStudy (anonymous):

but you have to divide the 160r^4 over Sqrt r^7? right?

hartnn (hartnn):

thats what i did....

OpenStudy (anonymous):

i know but like how do you divide it?

hartnn (hartnn):

now u just need to simplify... i will give a hint... \[\sqrt{a}=a^{1/2}\] and \[\frac{a^x}{a^y}=a^{x+y}\] try to solve it...if u can't i will explain afterwards...

OpenStudy (anonymous):

explain please ...... :(

hartnn (hartnn):

okk... \[\frac{r^4}{\sqrt{r^7}}=r^{4-7/2}=r^{1/2} = \sqrt{r}\] so final answer is \[160 \sqrt{r}\]

OpenStudy (anonymous):

um okay..thanks.really lost but yea .i am totally not ready for this complicated stuff in life. good luck becoming a professor :) you'll be great!

hartnn (hartnn):

thank u :)